-8x-3y=-4
7x+5y=13
x=
y=
I don't remember, if this is one of those substitution problems where you add or subtract, can you please help me figure out this problem.

Answer by stanbon(75887)

-8x-3y =-4
7x +5y =13
-----------------
Multiply thru the 1st equation by 5; Multiply thru the 2nd equation by 3:
--------------------------
-40x - 15y = -20
21x + 15y = 39
--------------------------
Add the two equations and solve for "x":
-19x = 19
x = -1
--------------
Substitute into -8x-3y=-4 to solve for "y":
-8*-1 -3y = -4
8 -3y = -4
-3y = -12
y = 4
==========
Cheers,
Stan H.